# Surface Reconstruction Using Rotation Systems

### v1

@@ 3.3. Triangulation
When edge insertion has completed, we initialize the output mesh, $M$ to be a copy of $T$. We then proceed to triangulate $M$ by greedily adding edges from $G$ to $M$ until no more triangles can be formed.
@@
The first step of triangulation is to compute the set of potential edges which could be added. We start by visiting all vertices of $M$, and for each vertex, $v$, we consider all outgoing halfedges. For each outgoing halfedge, $h_{1}=(v,u)$, we use the $\rho$ operator to find the next halfedge, $h_{2}=(v,w)$ and then the incident vertex, $w$, to which it points ($h_{1},h_{2}\in T$). Our edge candidate is now ${u,w}$. All edge candidates are pushed onto a priority queue, $q_{h}$, where the Euclidean length of prospective edges is the priority.
In each iteration of the triangulation loop, we pop an edge candidate to form a triangle if the edge passes the validity check which ensures both good topology and geometry. The validity check has three steps:
  • We check that the inserted edge (shown dashed in Figure 9) does not violate the RS. We check this by ensuring that the edges of the triangle are indeed adjacent in the cyclic ordering of all three vertices after the edge has been inserted. With reference to Figure 9 this means that $h_{1}=\rho(h_{2})\wedge h_{4}=\rho(h_{3})\wedge h_{6}=\rho(h_{5})$, where we also require that $\alpha<\pi\wedge\beta<\pi\wedge\gamma<\pi$. Without the angle constraint, flipped faces would be valid.
  • A heuristic face overlap check is performed to avoid the main cause of cracks in practice. For a triangle face, $f=(v,u,w)$, we project all vertices in the union of its $k$ nearest neighbors to the plane where the face lies, based on the face normal. If any of the vertices lies inside the triangle, the candidate is rejected. This approach serves as an effective method for preventing intersections and acts as a safeguard against undesirable configurations in which edge insertion is not feasible.
If the edge is valid, we update its halfedges with a new face identifier and insert its halfedges in the rotation system of $M$.
Figure 9. Illustration of RS inspections. The bottom row gives an invalid case. The dashed line corresponds to the edge considered for connection.
3.4. Reconstructing high genus meshes
As described above, this method can only reconstruct meshes of sphere or disk topology since no connections are made which could change the topology of the mesh. In order to facilitate reconstruction of objects with higher genus, we make a very simple adjustment to the algorithm.
As the pipeline in Figure 1 illustrates, for high-genus shape, two more operations called conservative triangulation and handle connection are required. The only difference between conservative triangulation and normal triangulation is that there is a threshold for edge length in conservative triangulation. This threshold is different for each vertex. It is implicitly encoded with the longest connection involving each vertex in the initial graph $G$. For instance, if a vertex $v$ has $k$ connected neighbors in $G$, the conservative edge length threshold of $v$ is determined by the length of the $2/3\cdot k$-th shortest edge associated with $v$. The reason for using a conservative triangulation step is that we thereby avoid skinny triangles encroaching on faces that are likely to be end-points of handles in the next step. Since a final triangulation stage follows handle insertion, we can afford to be conservative in the first step.
After undergoing conservative triangulation, the mesh $M$ attains a state where the majority of potential connections have been established. Assuming local edges share an average edge length $\overline{|e|}$, at this stage, an expected handle $\{u,w\}$ possesses a characteristic distinct from those of spurious handles: $d_{g}(u,w)\cdot\overline{|e|}\gg\overline{|e|}$, where $d_{g}$ is the graph distance. In practice, we use $d_{g}(u,w)>2n$ to detect if it is a handle we would like to connect.
To insert handles, we consider all edges of $G$ which have not been inserted in $M$ and whose end-points belong to different faces. If the edge fulfills the above condition, we enqueue it for insertion. Once all candidates have been considered, we insert the edges. The operation is exactly the same as the previous edge insertion except that when adding handles we only connect vertices which belong to different faces.
3.5. Dealing with noisy data
Scanned point clouds are always noisy since the points are based on measurements that are inherently subject to noise. Moreover, when we reconstruct from point clouds, we typically combine several sub-scans. These may not be perfectly aligned, introducing another source of noise.
Noise can be both tangential and normal to the surface. While tangential noise does not affect the reconstruction, normal noise can result in large Euclidean distances between points that are quite close in the tangent plane. To mitigate this issue, we use the estimated normals to project the points onto the tangent planes and compute the projected distances rather than the Euclidean distances.
To define projection distance, we need some preliminary definition: for edge $e=(v,u)$, $e_{//v}$ refers to the projection of $e$ on the tangential plane of $v$. The projection distance $|e|_{p}$ is defined
(4) $|e|_{p}=(|e_{//v}|+|e_{//u}|)/2\enspace.$
This approach enables the acquisition of a more noise resilient MST as the initial structure for reconstruction. However, an important consideration arises: when calling kd-tree in Euclidean distance at the initialization stage, it is conceivable for the initial graph $G$ to lack a valid connection between two vertices, even when the projection distance is small but the Euclidean distance is substantial due to noise. To mitigate this issue, a new kd-tree containing projected vertices is established and employed for all nearest-neighbor searches right from the outset.
4. Implementation
In this section, we discuss the practical aspects of the method in the order of the pipeline outlined in Section 3. These aspects include details of the quality test used in edge insertion, the binary tree data structure used to maintain faces, and the parameters used in the method.
4.1. Initialization
In the initialization phase, we construct a rotation system RS associated with the graph $G$. From a theoretical point of view, the RS is simply a cyclic ordering of neighboring vertices for each vertex of $G$, and the topology of the reconstructed surface is not dependent on which cyclic ordering we choose for each vertex. However, the geometry may become very convoluted if the cyclic orderings of two adjacent vertices are geometrically inconsistent.
Intuitively, geometric consistency means that if vertex $u$ is left of $v$ then $v$ must be right of $u$. Unfortunately, if we are dealing with a noisy point cloud and $u$ and $v$ are extremely close vertices in the sense of projection distance, small variations in the vertex normals can cause the cyclic orderings to be inconsistent. Typically in the example case depicted by Fig. 10, two vertices $u$ and $v$ possess an extremely short projection distance, a slight normal difference caused a situation that: both of them think they are at the top-left side of the other one. The neighbors of $u$ and $v$ are in the same order in the cyclic ordering of both vertices, e.g. given three neighbors $a$, $b$, and $c$, we have
    neighbors of $u$: $[a,b,c,v]$, and
    neighbors of $v$: $[a,b,c,u]$.
Since $u$ and $v$ have the same position in each other’s cyclic ordering, each is, effectively, top-left of the other. This is a problem from a geometric point of view since no straight edge can connect $u$ and $v$ consistently in this case. The situtation would lead to very convoluted faces.
In practice, the issue arises more often when the edge between two vertices is almost colinear with the normals of the vertices. Consequently, if the angle between the edge and the mean normal of the incident vertices is below a specified threshold $\phi$, the normal of one vertex is assigned to the other. This adjustment reduces the situation to the 2D case locally and ensures that there is a consistent connection between the two vertices.
Figure 10. The consequence of RS inconsistency. Note that the assumption here is that $u,v$ are extremely close to each other in 2D plane, this figure shows the rotation system of $u$. The rotation system will be the same for $v$ except for the swapped positions of $u$ and $v$. If the connection $\{u,v\}$ is allowed, convoluted faces will be created at the region of the orange shadow.
4.1.1. Analysis
@@
The initialization includes construction of a 3D kD-tree, computation of an MST, and initialization of the rotation system. Construction of such a kD-tree can be done in $O(N_{v}\log N_{v})$. With the kD-tree constructed, we generate a $k$NN graph by use of $k$NN queries on the tree. In the worst case, each query takes $O(N_{v})$ time to answer, but for practical purposes the expected running time of each query is $O(\log N_{v})$(Friedman et al., 1977). We can then compute the MST with an algorithm such as Prim’s algorithm in $O(kN_{v}\log N_{v})$ time. Finally we traverse the MST, constructing the rotation system by ordering of neighbours around each vertex.
@@
The worst case total time for initialization is thus $O(N_{v}^{2})$, due to the pessimistic bounds on answering $k$NN queries. For non-adversarial input, one would expect $O(kN_{v}\log N_{v})$ time.
4.2. Edge insertion
During the edge insertion stage, two tests are formulated to ascertain the soundness of both topology and geometry. Nevertheless, extensive experiments have revealed that the quality of connections established during this stage is pivotal in preventing holes, even though they are both topologically and geometrically closable. Consequently, a third practical test called the quality test, is implemented to uphold the connection’s high quality and align with intuitive expectations. Additionally, for enhanced efficiency, we have introduced a replacement setting for the geometry test.
4.2.1. Quality Test
In this test, several heuristics are employed to ensure the quality of the connections, as depicted in Fig. 11.
  • Firstly, an angle threshold, $\theta$, is imposed on the edge to which the leaf vertex $v$ is connected: $\alpha\geq\theta\wedge\beta\geq\theta$. This threshold ensures that we do not create extremely skinny triangles.
  • Next, we require that the normals of $v$ and the prospective connected vertex have a positive inner product, i.e. $\mathbf{n}_{v}\cdot\mathbf{n}_{u}>0$. This rule is needed to ensure that we do not make connections between opposite sides of thin structures.
  • Also, as we mentioned in Sec. 4.1, we want to avoid as many cases of inconsistency as possible. A solid inconsistency check is applied here: for any possible RS-valid triangles incident to this candidate, if the normals of three vertices point at the same side, it is consistent. Otherwise, any inconsistent case will get the candidate rejected.
  • Lastly, an angle threshold of $\gamma>90^{\circ}$ is imposed on the angle between the edge from $v$ to its parent vertex and the edge from parent to either grandparent or sibling. This helps ensure that we do not make connections to vertices that are too close in terms of graph distance. We avoid such connections since only a single connection is made from leaf vertices, and short connections are made in the triangulation phase.
Figure 11. An example configuration supporting the explanation of the quality test. The red dashed line is the connection to be considered.
4.2.2. Geometry test
In practice, considering efficiency, we use a fixed number $k/2$ of nearest neighbor vertices as the query set. At the same time, we pick a rejection set of $n\cdot k/2$ nearest neighbors to ensure that $r_{\mathrm{rej}}\gg r_{\mathrm{qry}}$. It turns out that this setting is more suitable considering both local point density and efficiency.
4.2.3. Analysis
@@
The running time of the edge insertion stage is dominated by the knn queries used to inspect the local geometry. Recall that we have constructed a kD-tree of the input points, and that a pessimistic worst case bound on answering these queries is $O(N_{v})$. Recall also that this edge insertion is applied to all leaves of the MST. It follows then that there are at most $O(N_{v})$ leaves to consider, each incurring $O(N_{v})$ worst case costs, totalling in $O(N_{v})^{2}$ time spent. In practice, the number of leaves might be much smaller, and the time to answer knn queries might behave closer to $O(\log N_{v})$.
4.3. Maintaining Faces
In order to efficiently perform the topology test described in Section 3.2.1, we need a data structure that allows us to quickly determine if two half-edges lie on the boundary of the same face, while also supporting fast updates as edges are inserted and faces split.
Using the Euler Tour technique (Henzinger and King, 1999), we store the sequence of half-edges of each face in a balanced binary tree with implicit keys. To query if two half-edges lie on the same face, we can simply traverse to the roots of the associated binary trees and compare, allowing for $O(\log N_{v})$ queries.
Storing the faces in this manner also allows for the use of standard binary tree operations, such as splits and joins, to be used in $O(\log N_{v})$ time. When inserting a new edge, we perform at most two splits to obtain binary trees representing the sequences of half-edges that bound the newly split face, and at most three joins to incorporate the new half-edges in these sequences. Thus, we can maintain the faces under insertion of new edges in $O(\log N_{v})$ time as well.
4.4. Triangulation
During triangulation, besides connecting edges, we need to generate the triangles incident to the connected edges. A common situation is that multiple possible triangles are incident to the candidate connection $\{u,v\}$. Fig. 12 illustrates a typical example where the red dashed line is the candidate connection. We iterate all the third vertex of the incident triangles other than $u$ and $v$. There will only be at most one triangle keeping a valid RS shown in Figure 9 for $h=(u,v)$. Similarly, for $h=(v,u)$, at most one triangle exists with valid RS. Thus, we can keep the 2-manifold of the output mesh and, at the same time, generate all valid triangles without leaving holes.
Figure 12. An example when a connection can trigger the generation of multiple triangles. The RS check will filter out all the invalid triangles and keep the 2-manifold property.
To perform triangulation, a bound on the worst case number of candidate edges is $O(kN_{v})$, since for each outgoing edge of each vertex we add exactly one candidate. We construct a priority queue on these candidate edges, perform our validity check, and potentially add them to our structures. The total cost per edge is then $O(\log N_{v})$, totaling in $O(kN_{v}\log N_{v})$ time to triangulate.
@@ 4.4. Triangulation
To check the distance criteria when reconstructing higher genus meshes, we run a restricted Dijkstra’s algorithm that stops early if the criteria has been met. Note that the number of edges in the graph is upper bounded by $O(kN_{v})$, since we are triangulating a subset of a $k$NN graph. Our worst case bound for this restricted search is thus $O(kN_{v}\log N_{v})$, attributed to the fact that we might not reach the early stopping condition, before the entire graph is explored.
Recall that the candidate edges for handle insertion are obtained from the $k$NN graph. There are thus $O(kN_{v})$ candidates in the worst case. Recall then that the criteria for insertion is that the distance along the graph between two endpoints of a candidate edge is longer than the euclidean distance. Thus, since adding edges to the graph cannot increase the distance along the graph, we need only examine each candidate once in order to either add it, or disqualify it from all future examinations.
Therefore we perform these restricted Dijkstra searches at most $O(kN_{v})$ times, totalling in $O(k^{2}N_{v}^{2}\log N_{v})$ time in the worst case.
As we run the algorithm for $k=30$, the total running time becomes $O(N_{v}^{2}\log N_{v})$ in the worst case.
@@ 4.5. Parameter settings
In Section 3 and Section 4, several parameters need to be set. Table  1 shows all the settings used in our implementation. While the settings were arrived at experimentally, the trade-offs between performance and efficiency are clear and explained in the following.
Variable name $k$ $n$ $\theta$ $\phi$
Value 30 5 30 10
Table 1. Practical parameter settings in the implementation.
The variable $k$ denotes the connectivity of the local neighborhood in the initial graph $G$. For optimal performance, creating a complete graph from the input point cloud is ideal, ensuring all potential edges are considered. However, achieving both time and memory efficiency in such a scenario is impractical. Therefore, under feasible conditions for running time and memory usage, a larger value of $k$ minimizes the likelihood of missing valid connections. Simultaneously, $n$ serves as a multiplier to establish a $\gg$ relationship; as $n$ increases, the result becomes more reliable. However, a trade-off exists, as larger values of $n$ lead to longer running times.
The parameters $\phi$ and $\theta$ represent angle thresholds for detecting inconsistency and measuring connection quality. An excessively large $\phi$ poses the risk of introducing new inconsistencies into the neighborhood, while an overly large $\theta$ can eliminate potentially good connections, causing consequences as illustrated in Figure 6.
@@ 5. Results
In this section, we first discuss an ablation study we conducted to provide an understanding of the tests in Section 3.2.
@@
To evaluate our method extensively, we carry out experiments on two different types of point cloud, specifically from synthetic data, i.e. data sampled from existing surfaces, and real scanning data. For the synthetic data sets, we assume that they are sampled sufficiently densely to admit an unambiguous reconstruction and that they are not strongly affected by noise.
@@
On the other hand, the real scans are often more dense since they tend to be composed of several combined scans. This makes it easy to estimate a relatively smooth normal field despite this type of data typically being affected by significant amounts of noise.
@@
We will compare our reconstructions to state-of-the-art combinatorial reconstruction methods and screened Poisson reconstruction. All the experiments were conducted on a laptop with a CPU 11th Gen Intel(R) Core(TM) i7-11800H @ 2.3GHz and 16 GB RAM.
5.1. Ablation study
As delineated in Section 3.2, the edge insertion stage involves three separate tests: the topology test $T_{t}$, the geometry test $T_{g}$, and the quality test $T_{q}$. To assess the importance of each test, we conducted an ablation study. In the baseline configuration all three tests are active. Generally, if we insert an edge that is undesirable, the consequence is that it prevents further insertion, leading to holes. Hence, the quantitative metric adopted is the count of boundary edges ($Bdry.e$). The raw range data of the Stanford Bunny is chosen as the test input due to its widespread recognition and its capacity to expose challenges encountered during edge insertion in our algorithm.
Figure 21 depicts the qualitative results, while Table 2 presents the quantitative results. Below, we analyze our findings and provide zoomed-in examples from each experimental setting.
Table 2. Statistics of ablation study. A larger number of boundary edges are left when any of the three tests are switched off, which means holes exist.
Experiment setting baseline w/o $T_{t}$ w/o $T_{g}$ w/o $T_{q}$
$Bdry.e$ 173 385 889 720
W/o topology test
Fig. 13 depicts a typical type of failure when the topology test is off. The red triangle in each picture connects two boundary edge loops. Effectively, this means that a spurious handle is introduced.
Figure 13. Failure cases when the topology test is switched off.
W/o geometry test
Fig. 14 illustrates a failure case occurring when the geometry test is disabled. All the holes remaining on the bunny exhibit a consistent type of configuration, referred to as a ”candy wrap” or ”twisted connection”. These are deemed highly undesirable from a geometry point of view, despite not being a topological problem. The geometry test serves to reject instances of the candy wraps.
a                                         b
Figure 14. Failure case when the geometry test is switched off. (a) shows part of a reconstructed mesh where the red edge is rejected by the geometry test. (b) shows the abstraction of this configuration from which we can see that the topology test would not reject the red edge, but if it is added we get faces (indicated by fat green and yellow lines) that are twisted in a fashion similar to candy wrapping.
W/o quality test
Fig. 15 demonstrates a failed example after switching off the quality test. If we construe the mesh and the surrounding hole as an ”island” and a ”lake” connected by a thin bridge, the issue is that due to a connection in the bridge made possible by disabling the quality check, the boundaries of the lake and the island are edge loops with different face ID’s. Hence, they will not be connected although they clearly belong to the same surface component.
Figure 15. Failure case when the quality test is switched off. The connection $\{u,w\}$ represented by the dashed line causes the boundary loop of mesh fragment inside the hole to become independent of the outer boundary loop. Hence, no connections are made between them.

### v2

@@ 4.3. Triangulation
After the handle edges have been inserted, the binary tree which stores the association of faces to halfedges is no longer valid. In principle, we could have kept it updated, but at this point, the remaining large faces (in terms of number of edges) correspond to holes, and the remaining small faces can be triangulated efficiently using the following procedure.
@@
The first step of triangulation is to compute the set of potential edges which could be added. We start by visiting all vertices associated with the connected handles. For each vertex, $u$, we consider all outgoing halfedges. For each outgoing halfedge, $h_{1}=(u,v)$, we use the $\rho$ operator to find the next halfedge, $h_{2}=(u,w)$ and then the incident vertex, $w$. Our edge candidate is now $\left\{v,w\right\}$. In most cases $\left\{v,w\right\}$ is already in $M$. Otherwise, adding the edge will ”clip” a triangle from the face, and it is pushed onto a priority queue, $Q_{t}$. In each iteration of the triangulation loop, we pop an edge candidate. The candidate edge is inserted if the corner of the opposite vertex (i.e. $u$) has not been split, since forming a triangle by inserting the edge would then lead to a non-manifold configuration.
If the edge is inserted, the incident vertices will be visited to search for more edge candidates. If no more candidates can be found, we traverse all vertices in $M$ that have not been visited to ensure all remaining faces are triangulated.
4.4. Time complexity analysis
4.4.1. Initialization
@@
The initialization includes the construction of a 3D kD-tree, computation of an MST, initialization of $RS$, and sorting edges in $G$. Construction of such a kD-tree can be done in $O(|V|\log|V|)$. With the kD-tree constructed, we generate $G$ by use of $k$NN queries on the tree. In the worst case, each query takes $O(|V|)$ time to answer, but for practical purposes, the expected running time of each query is $O(\log|V|)$(Friedman et al., 1977). We can then compute the MST with an algorithm such as Prim’s algorithm in $O(k|V|\log|V|)$ time. Finally, we traverse the MST and $G$ to construct $RS$ and sort the edges, where the worst case takes $O(|V|\log|V|)$.
@@
The worst-case total time for initialization is thus $O(|V|^{2})$, due to the pessimistic bounds on answering $k$NN queries. For non-adversarial input, one would expect $O(k|V|\log|V|)$ time.
4.4.2. Edge insertion
@@
The running time of the edge insertion stage is dominated by the $k$NN queries used to inspect the local geometry. Recall that we have constructed a kD-tree of the input points and that a pessimistic worst-case bound on answering these queries is $O(|V|)$. It follows then that there are at most $O(k|V|)$ edges to consider, each incurring $O(|V|)$ worst-case costs, totaling in $O(|V|^{2})$ time spent. In practice, the number of edges must be much smaller, and the time to answer $k$NN queries might behave closer to $O(\log|V|)$.
4.4.3. Handle connection
Recall that the shortest path between the incident vertices of each handle needs to be calculated. To do this, we run a restricted Dijkstra’s algorithm that stops early if the distance threshold has been reached.
@@ 4.4.3. Handle connection
Note that the number of edges in the graph is bounded from above by $O(k|V|)$ since we are searching through a subgraph of $G$. Our worst case bound for this restricted search is thus $O(k|V|\log|V|)$, attributed to the fact that we might not reach the early stopping condition before the entire graph is explored.
The candidate edges for handle insertion are obtained from $G$. There are thus $O(k|V|)$ candidates in the worst case. Therefore, we perform these restricted Dijkstra searches at most $O(k|V|)$ times, totaling in $O(k^{2}|V|^{2}\log|V|)$ time in the worst case, though this is highly unlikely to happen in practical usage.
4.4.4. Triangulation
To perform triangulation, a bound on the worst-case number of candidate edges is $O(k|V|)$. We construct a priority queue on these candidate edges, perform our validity check, and potentially add them to our structures. The total cost per edge is then $O(\log|V|)$, totaling in $O(k|V|\log|V|)$ time to triangulate.
4.4.5. Total time complexity
Taking the fact that the time complexity in the worst case of face identifier maintenance and triangulation is $O(\log|V|)$ and $O(k|V|\log|V|)$. Theoretically, the worst time complexity is dominated by handle connection, which is $O(k^{2}|V|^{2}\log|V|)$. In practice, however, the most time-consuming procedure is edge insertion since there are many more regular edges than handle edges even for objects of very high genus. In practice, we plot the time spent on the majority of tested shapes in Section 5, as depicted in Figure 11. The results demonstrate an almost linear increase in time as the input number of vertices, $|V|$, increases.
Figure 11. Time consumption on each shape in our results. Both axes are plotted in log scale, and the mesh name is annotated at the right of the data point. Notably, (s) stands for synthetic point cloud, while (r) stands for real scanning point cloud.
@@ 4.5. Parameter settings
Our method relies on several parameters. These are all discussed in Section 3. Below, we summarize the significance and default value of the four parameters we consider user exposable.
  • $k$ (default: $30$): the number of neighbors to which each point is connected in the initial graph $G$. It decides the initial connectivity of $G$. All edges connected in later stages are from this $k$NN graph.
  • $r$ (default: $20$): edges longer than $r$ times the average edge length are culled. Some vertices are outliers all of whose incident edges are significantly longer than the average edge length, and some connections are between points sampled from different surface components.
  • $\theta$ (default: $60^{\circ}$): edges in $G$ for which the angle between the normals associated with the endpoints is greater than $\theta$ are removed. $\theta$ is used to reject edges whose incident vertices’ normals are affected by significant noise, or edges connecting vertices from different surface components.
  • $n$ (default: $50$): if the number of edges in the shortest path between the endpoints of a handle edge is less than $n$ it is rejected. Local spurious handles are eliminated by setting this threshold.
It is reasonable to consider whether we should eschew the first three parameters and simply form a complete graph. In principle, this could be advantageous and would allow the algorithm to close holes of arbitrary size. Unfortunately, the cost in terms of both memory and run time would be prohibitive as indicated by the complexity analysis.
Figure 12. Renderings of models reconstructed using our method. Top image: well-known models from the Stanford 3D Scanning Repository. Middle image: several test objects are shown. The terrain (shown after hole closing) consists of two very close surfaces, the tree exhibits significant variation point density, and the two remaining objects are high-genus shapes. Bottom: we used our method to reconstruct the DTU STL0024 model as well as the Stanford Armadillo and Bunny directly from the acquired points. For each model, we see unprocessed reconstructions on the left and models after hole closing and denoising on the right. In all images, back-facing surfaces are rendered in red to show holes in the reconstruction.
@@ 5. Results
Table 1. Quantitative results on a diverse range of data. Models in the top half are from synthetic datasets while those in the bottom half are from scanning datasets. All four indicators are better when smaller. As most meshes are not originally watertight, the number of boundary edges will barely decrease to zero. To maintain a reasonable runtime, the point cloud from the DTU dataset (i.e. Stl002, Stl003, and Stl024) is reduced to approximately 500,000 points through random sampling.
Model Name $|V|$ Ours ScaleSpace Co3Ne w/o smoothing Co3Ne w/ smoothing DSE Meshing Poisson
$r_{v}(\%)\uparrow$ $|E_{b}|\downarrow$ $t(s)\downarrow$ $r_{v}(\%)\uparrow$ $|E_{b}|\downarrow$ $t(s)\downarrow$ $r_{v}(\%)\uparrow$ $|E_{b}|\downarrow$ $t(s)\downarrow$ $r_{v}(\%)\uparrow$ $|E_{b}|\downarrow$ $t(s)\downarrow$ $r_{v}(\%)\uparrow$ $|E_{b}|\downarrow$ $t(s)\downarrow$ —V— $t(s)\downarrow$
High-genus 2 288,182 99.996 200 46 52.926 100,850 97 99.982 3,720 0.5 99.769 11,465 0.3 100 783 764 253,730 7.7
Tree 982,923 99.998 2,574 482 63.125 580,363 3,725 99.893 23,581 1.2 99.936 28,851 1.3 - - - 1,770,349 56
Terrain 80,000 100 650 12 68.678 52,475 25 99.411 8,469 0.2 98.994 9,244 0.2 99.903 14,402 224 299,737 69
High-genus 1 143,662 100 0 21 99.965 0 19 100 338 0.2 100 534 0.2 100 106 391 336,222 9.0
Mobius band 76,387 100 0 10 26.897 20,656 186 99.999 386 0.1 98.606 14,291 0.2 100 12 187 104,756 3.4
Bunny 362,271 99.999 276 65 99.912 333 59 93.070 62,053 1.2 99.997 408 0.4 99.923 72,928 857 689,066 19
Armadillo 2,166,950 99.748 1,074 841 98.694 11,602 792 81.115 1,284,705 21.7 98.545 101,031 5.2 - - - 2,142,910 71
Stl002 499,501 99.999 455 137 99.868 2,268 644 90.890 177,184 3.7 99.987 701 0.9 99.968 76,443 1101 559,810 18
Stl003 498,768 99.997 1,879 121 97.073 5,244 714 90.129 158,215 3.5 99.916 2,401 0.6 99.904 106,544 1074 867,621 27
Stl024 500,000 99.999 1,387 116 98.316 2,171 831 94.914 101,693 2.8 99.971 2,853 0.4 99.977 52,972 1153 477,303 15
@@
To evaluate our method extensively, we carried out experiments on two different types of point clouds, namely synthetic data, i.e. data sampled from existing surfaces, and real scanning data.
@@
The synthetic data sets are typically not noisy, while real scans, on the other hand, are very dense since they tend to be composed of several combined sub-scans. This makes it easy to estimate a relatively smooth normal field despite this type of data being affected by significant amounts of noise.
@@
We will compare our reconstructions to several state-of-the-art combinatorial reconstruction methods as well as screened Poisson reconstruction. All experiments, except those that involve DSE Meshing, were conducted on a laptop with a CPU 13th Gen Intel(R) Core(TM) i7-1365U @ 1.8GHz and 32 GB RAM. As DSE Meshing requires GPU, it is tested on HPC with GPU Tesla V100-PCIE-16GB, CPU Intel(R) Xeon(R) CPU E5-2660 v3 @ 2.60GHz, and 128 GB RAM.